- The rate constant for the reaction is 0.330 M-.- at 200 °C. A — products If the initial concentration of A is 0,00120 M,
Answer #1


rate constant is k = 0.330 M -1 S-1

units of the K indicates that the order is second order

the integrated rate law is as follows

1 / [A] = kt + 1 / [A]0

here the [A] = concentration after time

[A]0 = initial concentration

k = rate constant = 0.330 M-1 s-1

t = time = 895 sec

substituting the values

1 / [A] = kt + 1 / [A]0

1 / [A] = 0.330 * 895 + 1 / 0.00120

1 / [A] = 0.330 * 895 + 833.33

1 / [A] = 1128.68

=> [A] = 0.0008859 M

answer => 0.0008859 M    (or) 8.86*10-4 M